# Recursive construction and enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic - I

### v1

@@ 1 Introduction
Let $GR(p^{\mathfrak{s}},m)$ denote the Galois ring of characteristic $p^{\mathfrak{s}}$ and cardinality $p^{\mathfrak{s}m},$ where $p$ is a prime and $m,$ $\mathfrak{s}$ are positive integers. By Theorem XVII.5 of [21], any finite commutative chain ring is isomorphic to a quotient ring of the form $\mathscr{R}_{e,m}=\frac{GR(p^{\mathfrak{s}},m)[y]}{\langle h(y),p^{\mathfrak{s}-1}y^{\mathtt{t}}\rangle},$ where $h(y)$ is an Eisenstein polynomial over $GR(p^{\mathfrak{s}},m)$ of degree $\kappa,$ and the parameter $\mathtt{t}$ satisfies $1\leq\mathtt{t}\leq\kappa$ when $\mathfrak{s}\geq 2$, while $\mathtt{t}=\kappa$ when $\mathfrak{s}=1.$ In the special case when $\mathfrak{s}=\kappa=\mathtt{t}=1,$ the ring $\mathscr{R}_{e,m}$ reduces to the finite field $\mathbb{F}_{p^{m}}$ of order $p^{m}$ and the enumeration formulae for self-orthogonal and self-dual codes over $\mathbb{F}_{p^{m}}$ are obtained by Pless [27]. Further, for a detailed study on the enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of odd characteristic (i.e., when $p$ is odd), the reader is referred to [4, 5, 14, 15, 23, 24, 33] and references therein.
@@ 2 Some preliminaries
A finite commutative ring with unity is called a chain ring if its ideals are totally ordered under the set-theoretic inclusion relation. The following theorem provides a complete characterization of all finite commutative chain rings.
Theorem 2.1.
[21] For a prime number $p$ and positive integers $m$ and $\mathfrak{s},$ let $GR(p^{\mathfrak{s}},m)$ denote the Galois ring of characteristic $p^{\mathfrak{s}}$ and cardinality $p^{\mathfrak{s}m}.$ The quotient ring $\mathcal{R}=\frac{GR(p^{\mathfrak{s}},m)[y]}{\langle h(y),p^{\mathfrak{s}-1}y^{\mathtt{t}}\rangle}$ is a finite commutative chain ring, where $h(y)$ is an Eisenstein polynomial over $GR(p^{\mathfrak{s}},m)$ of degree $\kappa$ and the parameter $\mathtt{t}$ satisfies $1\leq\mathtt{t}\leq\kappa$ when $\mathfrak{s}\geq 2$, while $\mathtt{t}=\kappa$ when $\mathfrak{s}=1.$ If $u:=y+\langle h(y),p^{\mathfrak{s}-1}y^{\mathtt{t}}\rangle\in\mathcal{R},$ then the ideal $\langle u\rangle$ is the unique maximal ideal of $\mathcal{R}$ of nilpotency index $e=\kappa(\mathfrak{s}-1)+\mathtt{t}$ and the residue field $\overline{\mathcal{R}}=\mathcal{R}/\langle u\rangle$ of $\mathcal{R}$ is of order $p^{m}.$ Moreover, the chain consisting of all ideals of $\mathcal{R}$ is given by $\{0\}\subsetneq\langle u^{e-1}\rangle\subsetneq\langle u^{e-2}\rangle\subsetneq\cdots\subsetneq\langle u\rangle\subsetneq\langle 1\rangle=\mathcal{R}$ and the element $p\in\langle u^{\kappa}\rangle\setminus\langle u^{\kappa+1}\rangle.$ Conversely, any finite commutative chain ring is isomorphic to a quotient ring of the form $\mathcal{R}$ for some prime number $p,$ positive integers $m$, $\mathfrak{s}$, $\kappa$ and $\mathtt{t}$ satisfying $1\leq\mathtt{t}\leq\kappa$ if $\mathfrak{s}\geq 2,$ while $\mathtt{t}=\kappa$ if $\mathfrak{s}=1,$ and for some Eisenstein polynomial $h(y)\in GR(p^{\mathfrak{s}},m)[y]$ of degree $\kappa,$ (note that the integers $p,$ $m,$ $\mathfrak{s},$ $\kappa$ and $\mathtt{t}$ are called invariants of the chain ring $\mathcal{R}$).
In particular, when $\mathfrak{s}=1,$ we see, by Theorem 2.1, that $\mathtt{t}=\kappa,$ $GR(p^{\mathfrak{s}},m)=\mathbb{F}_{p^{m}}$ and the ring $\mathcal{R}$ is the quasi-Galois ring $\mathbb{F}_{p^{m}}[y]/\langle y^{\kappa}\rangle,$ where $\mathbb{F}_{p^{m}}$ is the finite field of order $p^{m}.$ On the other hand, when $\kappa=1,$ we see, by Theorem 2.1 again, that $\mathtt{t}=1$ and the ring $\mathcal{R}$ is the Galois ring $GR(p^{\mathfrak{s}},m)$ of characteristic $p^{\mathfrak{s}}$ and cardinality $p^{\mathfrak{s}m}.$ Thus, quasi-Galois rings and Galois rings constitute two special classes of finite commutative chain rings.
@@
$\mathcal{G}=\begin{bmatrix}\mathtt{I}_{\lambda_{1}}&\mathtt{B}_{1,1}&\cdots&\mathtt{B}_{1,e-2}&\mathtt{B}_{1,e-1}&\mathtt{B}_{1,e}\\ 0&u\mathtt{I}_{\lambda_{2}}&\cdots&u\mathtt{B}_{2,e-2}&u\mathtt{B}_{2,e-1}&u\mathtt{B}_{2,e}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&u^{e-2}\mathtt{I}_{\lambda_{e-1}}&u^{e-2}\mathtt{B}_{e-1,e-1}&u^{e-2}\mathtt{B}_{e-1,e}\\ 0&0&\cdots&0&u^{e-1}\mathtt{I}_{\lambda_{e}}&u^{e-1}\mathtt{B}_{e,e}\end{bmatrix}=\begin{bmatrix}\mathtt{W}_{1}\\ u\mathtt{W}_{2}\\ \vdots\\ u^{e-2}\mathtt{W}_{e-1}\\ u^{e-1}\mathtt{W}_{e}\end{bmatrix}$ (2.1)
whose columns are partitioned into blocks of sizes $\lambda_{1}$, $\lambda_{2}$, $\ldots$, $\lambda_{e},$ $\lambda_{e+1}=n-(\lambda_{1}+\lambda_{2}+\cdots+\lambda_{e}),$ the matrix $\mathtt{I}_{\lambda_{i}}$ is the $\lambda_{i}\times\ \lambda_{i}$ identity matrix over $\mathscr{R}_{e,m}$ and the matrix $\mathtt{B}_{i,j}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathscr{R}_{e,m})$ is considered modulo $u^{j-i+1}$ for $1\leq i\leq j\leq e,$ i.e., the matrix $\mathtt{B}_{i,j}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathscr{R}_{e,m})$ is of the form $\mathtt{B}_{i,j}=\mathtt{B}_{i,j}^{(0)}+u\mathtt{B}_{i,j}^{(1)}+\cdots+u^{j-i}\mathtt{B}_{i,j}^{(j-i)}$ with $\mathtt{B}_{i,j}^{(0)},\mathtt{B}_{i,j}^{(1)},\ldots,\mathtt{B}_{i,j}^{(j-i)}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathcal{T}_{e,m})$ for $1\leq i\leq j\leq e,$ (throughout this paper, the set of all $f\times h$ matrices over a ring $R$ is denoted by $\mathcal{M}_{f\times h}(R)$). A linear code $\mathscr{D}$ of length $n$ over $\mathscr{R}_{e,m}$ is said to be of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ if it is permutation equivalent to a code with a generator matrix in the standard form (2.1). Further, a linear code $\mathscr{D}$ of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ over $\mathscr{R}_{e,m}$ contains $(p^{m})^{\sum\limits_{i=1}^{e}(e-i+1)\lambda_{i}}$ codewords.
@@
The set $\mathcal{T}_{m}^{n}$ consisting of all $n$-tuples over $\mathcal{T}_{m}$ can be viewed as an $n$-dimensional vector space over the finite field $\mathcal{T}_{m}$ under the component-wise addition induced by $\oplus$ and the component-wise scalar multiplication. We further define a map $B_{m}:\mathcal{T}_{m}^{n}\times\mathcal{T}_{m}^{n}\rightarrow\mathcal{T}_{m}$ as $B_{m}(\textbf{v},\textbf{w})=\pi_{0}(\textbf{v}\cdot\textbf{w})=\pi_{0}(\sum\limits_{i=1}^{n}v_{i}w_{i})$ for all $\textbf{v}=(v_{1},v_{2},\ldots,v_{n}),$ $\textbf{w}=(w_{1},w_{2},\ldots,w_{n})\in\mathcal{T}_{m}^{n},$ (here $\textbf{v}\cdot\textbf{w}=\sum\limits_{i=1}^{n}v_{i}w_{i}$ is viewed as an element of $\mathscr{R}_{e,m}$). Note that the map $B_{m}$ is a symmetric and non-degenerate bilinear form on $\mathcal{T}_{m}^{n}.$ A linear code $\mathcal{C}$ of length $n$ over $\mathcal{T}_{m}$ is defined as a $\mathcal{T}_{m}$-linear subspace of $\mathcal{T}_{m}^{n}.$ The dual code of $\mathcal{C}$ with respect to $B_{m}$ is defined as $\mathcal{C}^{\perp_{B_{m}}}=\{\textbf{w}\in\mathcal{T}_{m}^{n}:B_{m}(\textbf{w},\textbf{c})=0\text{ for all }\textbf{c}\in\mathcal{C}\}.$ Note that $\mathcal{C}^{\perp_{B_{m}}}$ is also a linear code of length $n$ over $\mathcal{T}_{m}.$ Further, a linear code $\mathcal{C}$ of length $n$ over $\mathcal{T}_{m}$ is said to be (i) self-orthogonal if it satisfies $\mathcal{C}\subseteq\mathcal{C}^{\perp_{B_{m}}},$ and (ii) self-dual if it satisfies $\mathcal{C}=\mathcal{C}^{\perp_{B_{m}}}.$

### v2

@@ 1 Introduction
Let $GR(p^{\mathfrak{s}},m)$ denote the Galois ring of characteristic $p^{\mathfrak{s}}$ and cardinality $p^{\mathfrak{s}m},$ where $p$ is a prime and $m,$ $\mathfrak{s}$ are positive integers. By Theorem XVII.5 of [24], any finite commutative chain ring is isomorphic to a quotient ring of the form
$\mathscr{R}_{e,m}=\frac{GR(p^{\mathfrak{s}},m)[y]}{\langle h(y),p^{\mathfrak{s}-1}y^{\mathtt{t}}\rangle},\vskip-2.84526pt$ (1.1)
where $h(y)$ is an Eisenstein polynomial over $GR(p^{\mathfrak{s}},m)$ of degree $\kappa,$ and the parameter $\mathtt{t}$ satisfies $1\leq\mathtt{t}\leq\kappa$ when $\mathfrak{s}\geq 2$, while $\mathtt{t}=\kappa$ when $\mathfrak{s}=1.$ The integers $p,$ $m,$ $\mathfrak{s},$ $\kappa$ and $\mathtt{t}$ are called invariants of the chain ring $\mathscr{R}_{e,m}.$
In the special case when $\mathfrak{s}=\kappa=\mathtt{t}=1,$ the ring $\mathscr{R}_{e,m}$ reduces to the finite field $\mathbb{F}_{p^{m}}$ of order $p^{m},$ and the enumeration formulae for self-orthogonal and self-dual codes over $\mathbb{F}_{p^{m}}$ are obtained by Pless [30]. Further, for a detailed study on the enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of odd characteristic (i.e., when $p$ is odd), the reader is referred to [4, 5, 15, 17, 26, 27, 36] and references therein.
On the other hand, when $p=2,$ the following important special cases have been considered in a series of papers:
  • If $m=1,$ $\mathfrak{s}\geq 1$ and $\kappa=1,$ then we have $\mathtt{t}=1$ and $\mathscr{R}_{e,m}\simeq\mathbb{Z}_{2^{\mathfrak{s}}}.$ Enumeration formulae for self-dual codes over $\mathbb{Z}_{2^{\mathfrak{s}}}$ are obtained by Nagata et al. [25].
  • If $\mathfrak{s}=1,$ then we have $\mathtt{t}=\kappa,$ $GR(2^{\mathfrak{s}},m)=\mathbb{F}_{2^{m}}$ and $\mathscr{R}_{e,m}\simeq\mathbb{F}_{2^{m}}[y]/\langle y^{\kappa}\rangle.$ Yadav and Sharma [35] provided a recursive method to construct and enumerate self-orthogonal and self-dual codes over $\mathbb{F}_{2^{m}}[y]/\langle y^{\kappa}\rangle$ from self-orthogonal codes over $\mathbb{F}_{2^{m}}.$
@@
    However, working as in Section 33 of [35], we observe that even these modified recursive methods are insufficient to construct and enumerate self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic in the case when $p=2$ and both $\kappa,s\geq 2,$ as illustrated in Example 2.1.
The main goal of this paper is to provide a modified recursive method for constructing and enumerating all self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic (i.e. $p=2$) when $\kappa\geq 3$ is odd. The complementary case, where $\kappa$ is even, will be addressed in a subsequent work [33], in which we develop a recursive method for constructing self-orthogonal and self-dual codes over $\mathscr{R}_{e,m}$ from a chain of self-orthogonal codes over finite fields that meet certain additional conditions.
Together, the aforementioned works provide a complete solution to the problem of enumeration of self-orthogonal and self-dual codes over finite commutative chain rings. Notably, as shown in [21, 23], the resulting enumeration formulae have broader applications, including the enumeration of self-orthogonal and self-dual quasi-abelian codes and self-orthogonal and self-dual Galois additive cyclic codes over finite commutative chain rings, as well as various other special classes of linear codes that decompose into direct sums of self-orthogonal and self-dual linear codes over finite commutative chain rings. Additionally, these enumeration formulae are instrumental in classifying such codes up to monomial equivalence [5, 10, 17, 34, 36]. It is worth noting that all monomially equivalent linear codes over finite commutative chain rings have the same homogenous and overweight distances. Consequently, these enumeration formulae are also instrumental in the effective implementation of search algorithms aimed at identifying new codes with optimal parameters [18, 29].
@@
$\mathcal{C}^{(1)}\subseteq\mathcal{C}^{(2)}\subseteq\cdots\subseteq\mathcal{C}^{(\lceil\frac{e}{2}\rceil)}\vskip-5.69054pt$
of self-orthogonal codes of length $n$ over $\mathcal{T}_{m},$ and vice versa, under the following conditions:
  1. (i)
  2. (ii)
  3. (iii)
where $\lambda_{1},\lambda_{2},\ldots,\lambda_{e}$ are non-negative integers satisfying $2\lambda_{1}+2\lambda_{2}+\cdots+2\lambda_{e-i+1}+\lambda_{e-i+2}+\lambda_{e-i+3}+\cdots+\lambda_{i}\leq n$ for $\lceil\frac{e+1}{2}\rceil\leq i\leq e$ (see Propositions 3.13.5 and Figures 1 and 2). In Section 4, we employ the recursive construction methods outlined in Figures 1 and 2 to derive explicit enumeration formulae for all self-orthogonal and self-dual codes of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ (Theorems 4.1 and 4.2). Using Remark 2.1 and equation (4.1), these results further yield enumeration formulae for all self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m}.$ It is worth noting that when $\mathfrak{s}=1,$ we have $e=\mathtt{t}=\kappa,$ and the ring $\mathscr{R}_{e,m}$ reduces to the quasi-Galois ring $\mathbb{F}_{2^{m}}[y]/\langle y^{\kappa}\rangle.$ In this special case, for any odd $\kappa\geq 3,$ the enumeration formulae for self-orthogonal and self-dual codes over $\mathbb{F}_{2^{m}}\text{[}y\text{]}/\langle y^{\kappa}\rangle$ can be directly obtained from Theorems 4.1 and 4.2 upon setting $\mathfrak{s}=1$ and $e=\mathtt{t}=\kappa$ (see equation (4.1) and Remarks 4.2 and 4.3).
@@ 2 Some preliminaries
In this section, we will present fundamental properties of finite commutative chain rings and outline the theory of linear codes over such rings, including their generator matrices, duality and torsion codes. We will also explain the structural properties of self-orthogonal and self-dual codes and highlight the limitations of existing recursive constructions for these codes over finite commutative chain rings of even characteristic that are neither quasi-Galois nor Galois rings.
2.1 Finite commutative chain rings
A finite commutative ring with unity is called a chain ring if its ideals are totally ordered under the set-theoretic inclusion relation. Throughout this paper, let $p$ be a prime number, and let $m,$ $\mathfrak{s},$ $\kappa$ and $\mathtt{t}$ be positive integers satisfying $1\leq\mathtt{t}\leq\kappa$ if $\mathfrak{s}\geq 2,$ whereas $\mathtt{t}=\kappa$ if $\mathfrak{s}=1.$ Let $\mathscr{R}_{e,m}$ (as defined in (1.1)) be a finite commutative chain ring with invariants $p,$ $m,$ $\mathfrak{s},$ $\kappa,$ $\mathtt{t}$ and maximal ideal $\langle u\rangle$ of nilpotency index $e=\kappa(\mathfrak{s}-1)+\mathtt{t}.$ Here, the residue field $\overline{\mathscr{R}}_{e,m}=\mathscr{R}_{e,m}/\langle u\rangle$ is the finite field of order $p^{m}.$ By Theorem XVII.5 of [24], we note that all ideals of $\mathscr{R}_{e,m}$ form a chain $\{0\}\subsetneq\langle u^{e-1}\rangle\subsetneq\langle u^{e-2}\rangle\subsetneq\ldots\subsetneq\langle u\rangle\subsetneq\langle 1\rangle=\mathscr{R}_{e,m}$ and that the element $p\in\langle u^{\kappa}\rangle\setminus\langle u^{\kappa+1}\rangle.$ We further note, by Lemma XVII.4 of [24], that $|\langle u^{i}\rangle|=p^{m(e-i)}$ for $0\leq i\leq e,$ (throughout this paper, $|\cdot|$ denotes the cardinality function). By Theorem XVII.5 of [24] again, we see that there exists an element $\xi\in\mathscr{R}_{e,m}$ with multiplicative order $p^{m}-1$. Moreover, the cyclic subgroup generated by $\xi$ is the only subgroup of the unit group of $\mathscr{R}_{e,m}$, which is isomorphic to the multiplicative group of the residue field $\overline{\mathscr{R}}_{e,m}.$ The set
$\mathcal{T}_{e,m}=\{0,1,\xi,\xi^{2},\ldots,\xi^{p^{m}-2}\}\vskip-5.69054pt$
@@ 2.2 Linear codes over $\mathscr{R}_{e,m}$
$\mathcal{G}=\begin{bmatrix}\mathtt{I}_{\lambda_{1}}&\mathtt{B}_{1,1}&\cdots&\mathtt{B}_{1,e-2}&\mathtt{B}_{1,e-1}&\mathtt{B}_{1,e}\\ 0&u\mathtt{I}_{\lambda_{2}}&\cdots&u\mathtt{B}_{2,e-2}&u\mathtt{B}_{2,e-1}&u\mathtt{B}_{2,e}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&u^{e-2}\mathtt{I}_{\lambda_{e-1}}&u^{e-2}\mathtt{B}_{e-1,e-1}&u^{e-2}\mathtt{B}_{e-1,e}\\ 0&0&\cdots&0&u^{e-1}\mathtt{I}_{\lambda_{e}}&u^{e-1}\mathtt{B}_{e,e}\end{bmatrix}=\begin{bmatrix}\mathtt{W}_{1}\\ u\mathtt{W}_{2}\\ \vdots\\ u^{e-2}\mathtt{W}_{e-1}\\ u^{e-1}\mathtt{W}_{e}\end{bmatrix}$, (2.1)
where the columns of $\mathcal{G}$ are partitioned into blocks of sizes $\lambda_{1}$, $\lambda_{2}$, $\ldots$, $\lambda_{e},$ $\lambda_{e+1}=n-(\lambda_{1}+\lambda_{2}+\cdots+\lambda_{e}),$ $\mathtt{I}_{\lambda_{i}}$ denotes the $\lambda_{i}\times\ \lambda_{i}$ identity matrix over $\mathscr{R}_{e,m},$ and $\mathtt{B}_{i,j}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathscr{R}_{e,m})$ is considered modulo $u^{j-i+1},$ i.e., $\mathtt{B}_{i,j}=\mathtt{B}_{i,j}^{(0)}+u\mathtt{B}_{i,j}^{(1)}+\cdots+u^{j-i}\mathtt{B}_{i,j}^{(j-i)}$ with $\mathtt{B}_{i,j}^{(0)},\mathtt{B}_{i,j}^{(1)},\ldots,\mathtt{B}_{i,j}^{(j-i)}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathcal{T}_{e,m})$ for $1\leq i\leq j\leq e,$ (throughout this paper, the set of all $f\times h$ matrices over a ring $R$ is denoted by $\mathcal{M}_{f\times h}(R)$). A linear code $\mathscr{D}$ of length $n$ over $\mathscr{R}_{e,m}$ is said to be of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ if it is permutation equivalent to a code with a generator matrix in the standard form (2.1). Further, a linear code $\mathscr{D}$ of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ over $\mathscr{R}_{e,m}$ contains $(p^{m})^{\sum\limits_{i=1}^{e}(e-i+1)\lambda_{i}}$ codewords.
@@
2.4 Torsion codes over $\mathcal{T}_{m}$ and self-orthogonal and self-dual codes over $\mathscr{R}_{e,m}$
The set $\mathcal{T}_{m}^{n},$ consisting of all $n$-tuples over $\mathcal{T}_{m}$, can be viewed as an $n$-dimensional vector space over the finite field $\mathcal{T}_{m}$ under the component-wise addition induced by $\oplus$ and the component-wise scalar multiplication. We further define a map $B_{m}:\mathcal{T}_{m}^{n}\times\mathcal{T}_{m}^{n}\rightarrow\mathcal{T}_{m}$ as $B_{m}(\textbf{v},\textbf{w})=\pi_{0}(\textbf{v}\cdot\textbf{w})=\pi_{0}(\sum\limits_{i=1}^{n}v_{i}w_{i})$ for all $\textbf{v}=(v_{1},v_{2},\ldots,v_{n}),$ $\textbf{w}=(w_{1},w_{2},\ldots,w_{n})\in\mathcal{T}_{m}^{n},$ (here $\textbf{v}\cdot\textbf{w}=\sum\limits_{i=1}^{n}v_{i}w_{i}$ is viewed as an element of $\mathscr{R}_{e,m}$). Note that the map $B_{m}$ is a symmetric and non-degenerate bilinear form on $\mathcal{T}_{m}^{n}.$ A linear code $\mathcal{C}$ of length $n$ over $\mathcal{T}_{m}$ is defined as a $\mathcal{T}_{m}$-linear subspace of $\mathcal{T}_{m}^{n}.$ The dual code of $\mathcal{C}$ with respect to $B_{m}$ is defined as $\mathcal{C}^{\perp_{B_{m}}}=\{\textbf{w}\in\mathcal{T}_{m}^{n}:B_{m}(\textbf{w},\textbf{c})=0\text{ for all }\textbf{c}\in\mathcal{C}\}.$ Note that $\mathcal{C}^{\perp_{B_{m}}}$ is also a linear code of length $n$ over $\mathcal{T}_{m}.$ Further, a linear code $\mathcal{C}$ of length $n$ over $\mathcal{T}_{m}$ is said to be (i) self-orthogonal if it satisfies $\mathcal{C}\subseteq\mathcal{C}^{\perp_{B_{m}}},$ and (ii) self-dual if it satisfies $\mathcal{C}=\mathcal{C}^{\perp_{B_{m}}}.$